OpenStudy (anonymous):

Solve by the substitution method 8x+9y=-5 x=8-4y Answer should be in an ordered pair.

OpenStudy (lalaly):

substitute x=8-4y in 8x+9y=-5 8(8-4y)+9y=-5 64-32y+9y=-5 -23y=-69 y=3 substitute y=3 in sec equation x=8-12=-4

OpenStudy (lalaly):

it said substitution method not eliminaation

OpenStudy (anonymous):

lalaly thank you when I substitute 3 in the second equation will you double check it for me?

OpenStudy (lalaly):

its right... x=-4 and y=3 ;)

OpenStudy (anonymous):

OH ...LOL see I am not very good at algebra...haha I thought I had to go back and solve some more.. Thank you! If only I could give double medals.

OpenStudy (lalaly):

hehe im glad i could help :D

OpenStudy (anonymous):

Thank you again!